OpenStudy (anonymous):

Find a scalar S such that ||v+Sw|| = 6 when v=<-2,5> and w=<3,-2>

OpenStudy (anonymous):

|| <num1, num2, ...> || means sqrt( num1^2 + num2^2 + ...) <-2, 5> + S<3, -2> =<-2, 5> + <3S, -2S> =<-2 + 3S, 5 -2S> || <-2 + 3S, 5 -2S> || = sqrt( (-2+3S)^2 + (5-2S)^2 ) We want sqrt( (-2+3S)^2 + (5-2S)^2 ) to equal 6. Following so far?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Are you able to finish it?

OpenStudy (anonymous):

No this is where I got stuck haha

OpenStudy (anonymous):

Try squaring both sides. Next you'll expand the squares. And then just solve for S.

OpenStudy (zzr0ck3r):

@FutureMathProfessor are you ok? you just asked how to solve sqrt(something) = something?

OpenStudy (anonymous):

Matlab gives me: 347^(1/2)/13 + 16/13 and 16/13 - 347^(1/2)/13 :S Maybe this is why it's so difficult? Let me check this by hand...

OpenStudy (anonymous):

You end up with sqrt(13x^2 - 32x + 29) = 6 which definately doesn't come out to anything nice...